OpenStudy (anonymous):

It took John 12 hours riding his bike to make the round trip to his uncle's. If he averaged 20 mph out and 30 mph back, how long did he travel each way? (Round answer to nearest tenth.)

OpenStudy (anonymous):

if d is distace...then time taken t1=d/20 and t2= d/30 now t1+t2=12 d/20+d/30 = 12 d(1/20 +1/30)=12 d=144 m

OpenStudy (anonymous):

The distance remains the same. (12 - A) * 20 = (12 - B) * 30 The total time remains the same: A + B = 12 http://www.wolframalpha.com/input/?i=%2812-A%29*20+%3D+%2812-B%29*30+%26+A+%2B+B+%3D+12

OpenStudy (anonymous):

ok how many hours did he go each way?

OpenStudy (anonymous):

7.2 and 4.8

OpenStudy (anonymous):

144/20 =7.2 144/30=4.8